given n(E')=x and P(E)=y/(x+y), find n(S), n(E), P(E), odds (E).

Please also explain the methods behind doing so. Thank you so much!!!
Answer #1

P(E) = y/(x+y)

P(E') = 1-P(E) = x/(x+y)

as n(E') = x hence n(S) = x+y

P(E) = n(E)/n(S) hence n(E) = y

P(E)=y/(x+y)

P(E') = x/(x+y)

odds (E) is y:x

press like